MPC=0.4
Size of imports and taxes =0
What will be the size of aggregate expenditure multiplier?
1
Expert's answer
2011-12-06T10:58:27-0500
Aggregate expenditure multiplier = 1 / MPS, MPS = 1 - MPC, so we have:

aggregate expenditure multiplier = 1/ (1 - 0.4) = 5/3 ≈ 1.67